OpenStudy (anonymous):

complete the square on the quadratic state the vertex, min or max values, axis of symmetry y=-2x^2+3x+6

OpenStudy (anonymous):

\[y=-2(x-\frac{3}{4})^2+\text { something}\] and we can find the something easy enough

OpenStudy (anonymous):

since \[(x-\frac{3}{4})^2=0\iff x=\frac{3}{4}\] replace x by 3/4 in your original expression to find the something. then because this si a parabola that opens down, it will be a maximum.

OpenStudy (anonymous):

i believe it is \[\frac{57}{8}\] but you should check my arithmetic

OpenStudy (anonymous):

this is the max correct? how do I find the square.

OpenStudy (anonymous):

you find the vertex via \[x=-\frac{b}{2a}\] in this case you have \[a=-2,b=3,-\frac{b}{2a}=\frac{3}{4}\] so you know it will look like \[y=-2(x-\frac{3}{4})^2+k\] and you find k by substitution

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

there is a harder way to do this, but if you can remember \[-\frac{b}{2a}\] (which is not so bad) then you will be in good shape

OpenStudy (anonymous):

ok thanks